Sir Here is my question, Q. If p and q are the roots of the equation ax 2 +bx+c=0 & r and -q are the roots of lx 2 +mx+n=0,then show that p and r are the roots of (b/c+m/n)x 2 + (b/a+m/l)(b/c+m/n)x + (b/a+m/l) = 0 Hope to get an answer soon.

Sir


Here is my question,


Q. If p and  q are the roots of the equation ax2+bx+c=0 & r and -q are the roots of lx2+mx+n=0,then show that p and r are the roots of  (b/c+m/n)x2 + (b/a+m/l)(b/c+m/n)x + (b/a+m/l) = 0


Hope to get an answer soon.

Grade:12

1 Answers

Askiitian.Expert Rajat
14 years ago

p+q=-b/a              pq=c/a

r-q=-m/l                -qr=n/l

Let p,r be roots of

tx2+vx+w=0

then p+r = -v/t

and pr=w/t

As you know, p+r = -b/a - m/l

pr = ?

Now Solve these two equations for q2:

aq2+bq+c=0-------------------X m ................1

lq2-mq+n=0...................X b.................2

 

Add 1 and 2

q2 = - [mc+nb]/[am+lb]

-prq2= cn/al

Hence find out

pr = [m/l + b/a]/[m/n+b/c]=

p+r = -[b/a + m/l]

Also

p+r = -v/t

pr=w/t


Subtitue values for

w/t = [m/l + b/a]/[m/n+b/c]

and

-v/t = -[b/a + m/l]

And finally you'll get the required equation.